### v1

# Stochastic Monkeys at Play: Random Augmentations Cheaply Break LLM Safety Alignment

Abstract
Content warning: This paper contains examples of harmful language.
Safety alignment of Large Language Models (LLMs) has recently become a critical objective of model developers. In response, a growing body of work has been investigating how safety alignment can be bypassed through various jailbreaking methods, such as adversarial attacks. However, these jailbreak methods can be rather costly or involve a non-trivial amount of creativity and effort, introducing the assumption that malicious users are high-resource or sophisticated. In this paper, we study how simple random augmentations to the input prompt affect safety alignment effectiveness in state-of-the-art LLMs, such as Llama 3 and Qwen 2. We perform an in-depth evaluation of 17 different models and investigate the intersection of safety under random augmentations with multiple dimensions: augmentation type, model size, quantization, fine-tuning-based defenses, and decoding strategies (e.g., sampling temperature). We show that low-resource and unsophisticated attackers, i.e. stochastic monkeys, can significantly improve their chances of bypassing alignment with just 25 random augmentations per prompt.
1 Introduction
Autoregressive Large Language Models (LLMs) have become increasingly ubiquitous in recent years. A primary driving force behind the explosion in popularity of LLMs has been their application to conversational AI; e.g., chatbots that can engage in turn-by-turn conversation with humans (OpenAI, 2022). However, as the capabilities of LLMs have increased over the years, so have concerns about their potential for misuse by malicious users. In response to these concerns, tremendous efforts have been invested towards aligning LLMs (Ouyang et al., 2022; Rafailov et al., 2024; Ethayarajh et al., 2024). In order to safety-align a model, an extensive amount of manually-labeled preference data may be required to achieve a high quality alignment. Given the extensive investments required to align a model, it is critical for model developers to ensure that the alignment can withstand a broad range of real-world behavior from malicious users.
Unfortunately, it has been shown that safety alignment can be bypassed through a variety of techniques. One popular set of techniques is jailbreaks, where a malicious user modifies a harmful prompt in such a way that the aligned model complies with the request. These jailbreaks can either be manually crafted through clever prompt engineering (Liu et al., 2023), or automatically discovered using optimization-based adversarial attacks (Zou et al., 2023). In the former case, a nontrivial amount of creativity and effort may be required to create effective jailbreaks. In the latter case, only malicious users that have access to sufficiently powerful hardware may leverage such attacks. As such, one may wonder whether there are any simpler ways to effectively bypass safety alignment.
A recent number of works have shown that it is indeed possible to circumvent safety alignment with much less sophisticated methods (Huang et al., 2023; Andriushchenko & Flammarion, 2024; Vega et al., 2023). Such methods showcase how techniques to bypass safety alignment sits on a wide spectrum of complexity, with adversarial attacks occupying the high end of this spectrum. We hypothesize that effective random attacks, namely the simple use of random input augmentations, may exist on the low end of this spectrum. In the context of NLP, prior work investigating random augmentation attacks have largely focused on their impact to accuracy for classifier models  (Li et al., 2018; Morris et al., 2020; Zhang et al., 2021). Some recent work has started to explore their role in impacting safety for generative models, but only for purposes of defending the model  (Robey et al., 2023; Zhang et al., 2024). Hence, there is a critical gap to fill in evaluating their effectiveness for attacking generative model safety. (See Appendix A for more discussion on related work.)
Figure 1: An overview of the threat model we investigate. A malicious user (i.e. the stochastic monkey) randomly and independently augments the prompt $k$ times and observes $k$ different outputs. The attacker is successful if at least one of the outputs is compliant. Here, we show a successful example obtained from Llama 3.1 8B Instruct with $k=25$ using greedy decoding.
In this work, we address this gap by investigating a simple yet surprisingly under-explored question: how effectively can random augmentations bypass the safety alignment of state-of-the-art LLMs? In contrast to adversarial attacks, a simple application of random augmentations does not require any feedback from the model or intricate search processes, and is thus computationally cheap and algorithmically unsophisticated. As such, they can be easily utilized by a class of attackers we refer to as stochastic monkeys. Yet, despite their relative simplicity, we find that random augmentations can be surprisingly effective at eliciting compliant responses to harmful prompts. For instance, Figure 1 shows a real example where a compliant response was obtained from Llama 3.1 8B Instruct (Dubey et al., 2024) within just 25 augmentations that randomly changed just a single character.
Our key contributions and observations are as follows:
  1. 1.
    We investigate the effectiveness of simple character-level and string insertion random augmentations (see Table 1) towards bypassing safety alignment. We examine how safety under random augmentations is affected when varying the following aspects: augmentation type, model size, quantization, fine-tuning-based defenses, and decoding strategies (e.g., sampling temperature).
  2. 2.
[…]
  3. 3.
    We also observe that: 1 Character-level augmentations tend to be much more effective than string insertion augmentations for increasing success rate, 2 Larger models tend to be safer, 3 More aggressive weight quantization tends to be less safe, 4 Adversarial training can generalize to random augmentations, but its effect can be circumvented by decreasing augmentation intensity, and 5 Even when altering the sampling temperature, random augmentations still provide further improvements to the success rate. We also employ a human study on a sample of 1220 data points from our experiments to calibrate our evaluation metric for controlling the estimated false positive and false negative rates.
2 Evaluation Dimensions and Metric
2.1 Preliminaries
In this section, we introduce various notation and terminology used in our paper, as well as the primary aspects of our experiment pipeline.
Sequences and models. Let $V=\{1,2,\ldots,m\}$ represent a vocabulary of $m$ token, and let $\Sigma$ denote the set of printable ASCII characters. Let $\Sigma^{+}$ denote the set of positive-length sequences. An autoregressive LLM $f$ operates as follows: given an initial character sequence from $\Sigma^{+}$, $f$ outputs a probability distribution over $V$ to predict the next token (for simplicity, we view the tokenizer associated with $f$ as a part of $f$).
Generation. Model $f$ may be used as part of a broader pipeline where the input and output character sequences can be restricted to spaces $\mathcal{X}\subseteq\Sigma^{+}$ and $\mathcal{Y}\subseteq\Sigma^{+}$, respectively (e.g., with prompt templates, limits on sequence length, etc.). For simplicity, we define a generation algorithm $g$ to be this entire pipeline, which given $\textbf{x}\in\mathcal{X}$, uses $f$ to generate $\textbf{y}\in\mathcal{Y}$ following some decoding strategy. For generality, we assume $g$ to be stochastic, with deterministic algorithms being a special case.
Augmentations. An augmentation $a\colon\mathcal{X}\rightarrow\mathcal{X}$ is a function that modifies x before being passed to $g$. Note that “no augmentation” can be considered a special case where the “augmentation” is the identity function $a(\textbf{x})=\textbf{x}$. Let an augmentation set $\mathcal{A}$ be a set of augmentations that may be related in nature (e.g., appending a suffix of a specific length); we refer to the nature of the relation as the augmentation “type”. Augmentations may be randomly sampled, so we also associate a sampling distribution $P_{\text{aug}}(\,\cdot\,;\mathcal{A})$ with each $\mathcal{A}$. We let $\mathcal{A}_{I}$ denote the “no augmentation” singleton containing the identity function that is drawn with probability 1 from $P_{\text{aug}}(\,\cdot\,;\mathcal{A}_{I})$.
Safety dataset. For safety evaluation, we set $P_{\text{test}}$ to be an underlying distribution of inputs from $\mathcal{X}$ that contain harmful user requests. We assume that a finite set $\mathcal{D}$ of i.i.d. samples from $P_{\text{test}}$ is available. As what is deemed “harmful” is subjective and may change over time, we make no further assumptions about $P_{\text{test}}$ and simply assume that $\mathcal{D}$ is representative of the desired $P_{\text{test}}$.
Safety judge. A safety judge $c\colon\mathcal{X},\mathcal{Y}\rightarrow\{0,1\}$ outputs 1 if y is deemed compliant with a user request x and 0 otherwise. Different judges may involve different criteria for compliance. For simplicity, we assume part of $c$ includes any necessary preparation of x and y (e.g., removing the prompt template from x, applying a new prompt template, etc.). We always evaluate the compliance of y with respect to the original prompt, even if y was generated from an augmentation.
2.2 Research Questions
Our experiment pipeline has three main components that can be varied: the augmentation type, the model, and the generation algorithm. We will investigate how each of these components impact safety while isolating the other components, and therefore naturally split our research question into the following sub-questions:
RQ1. For a given model and generation algorithm, how do different augmentation types impact safety? There are many ways to randomly augment a prompt such that its semantic meaning is preserved (or at least highly inferable). However, there may be significant differences in how effectively they enable malicious users to bypass safety alignment. Hence, we examine how a variety of random augmentations can improve attack success over the baseline of not using any augmentations.
RQ2. For a given augmentation type and generation algorithm, how do different model aspects impact safety; specifically: model size, quantization and fine-tuning-based defense? Model developers commonly release models of multiple sizes within a model family, permitting accessibility to a broader range of hardware. Alternatively, extensive efforts have been made recently to quantize LLMs for similar reasons. Orthogonal to the goal of accessibility is how to make models safer against jailbreaks, for which some recent works have proposed fine-tuning-based defense methods. Hence, it is of practical interest to examine how the safety under random augmentations interacts with each of these aspects.
RQ3. For a given model, how much do random augmentations impact safety when different decoding strategies are used? By default, all our experiments are conducted using greedy decoding, so the no augmentation baseline in RQ1 only produces a single output per prompt. A critical question therefore is whether random augmentations provide any additional influence on success rates when $k$ random outputs are also sampled in the no augmentation case. Hence, we examine decoding strategies beyond greedy decoding.
2.3 Evaluation Metric
In realistic settings, a malicious user who seeks to elicit specific harmful content from an LLM may make multiple attempts before moving on. We therefore assume that for each harmful prompt $\textbf{x}_{i}\in\mathcal{X}$, a malicious user makes $k$ attempts where for each attempt a separate augmentation is first applied to the prompt, as illustrated in Figure 1. To evaluate success, we check whether the proportion of augmentations that produce outputs where safety judge $c$ evaluates to 1 is strictly greater than some threshold $\gamma\in[0,1)$. We refer to such an occurrence as a $(k,\gamma)$-success and define the following function for it:
$s_{k,\gamma}(\textbf{x},\textbf{y}_{1},\ldots,\textbf{y}_{k})\vcentcolon=\begin{cases}1&\text{if }\frac{1}{k}\sum\limits_{j=1}^{k}c(\textbf{x},\textbf{y}_{j})>\gamma\\ 0&\text{otherwise}\end{cases}$ (1)
where for $1\leq j\leq k$, $\textbf{y}_{j}\in\mathcal{Y}$ is the observed output given $a_{j}(\textbf{x})$, where $a_{j}\in\mathcal{A}$ is the $j$th observed augmentation. Note that the definition of $(k,\gamma)$-success has also been used as the majority vote definition for SmoothLLM  (Robey et al., 2023), although SmoothLLM uses Equation 1 solely as part of a defense mechanism whereas we use it for attack evaluation (see Appendix A.3).
Given we use a learned classifier for $c$, simply checking if any (i.e., $\gamma=0$) augmentation succeeds can have a high false positive rate (a false positive occurs when $s_{k,\gamma}(\textbf{x},\textbf{y}_{1},\ldots,\textbf{y}_{k})$ evaluates to 1 when in fact none of the $k$ outputs are harmful). A non-zero $\gamma$ can therefore be used to help reduce the false positive rate. However, applying too high of a threshold may result in a high false negative rate (a false negative occurs when $s_{k,\gamma}(\textbf{x},\textbf{y}_{1},\ldots,\textbf{y}_{k})$ evaluates to 0 when in fact at least one of the $k$ outputs are harmful). Thus, $\gamma$ should be carefully chosen so as to balance the false positive and false negative rates. See Appendix B.1 for more details.
Let $\textbf{X}_{\text{harm}}\sim P_{\text{test}}$ be a random harmful input prompt and $A_{1},\ldots A_{k},\stackrel{{\scriptstyle i.i.d.}}{{\sim}}P_{\text{aug}}(\,\cdot\,;\mathcal{A})$ be $k$ random augmentations from $\mathcal{A}$ to apply to $\textbf{X}_{\text{harm}}$ before being provided as $k$ inputs to $g$. Let $\textbf{Y}\,|\,\textbf{X}=\textbf{x}\sim P_{\textbf{Y}\,|\,\textbf{X}}(\,\cdot\,|\,\textbf{X}=\textbf{x};f,g)$ be a random output sequence from $\mathcal{Y}$ produced by $g$ using $f$, given an input $\textbf{x}\in\mathcal{X}$. Similarly, for $1\leq j\leq k$, let $\textbf{Y}_{j}\,|\,\textbf{X}_{\text{harm}}=\textbf{x},A_{j}=a_{j}\sim P_{\textbf{Y}\,|\,\textbf{X}}(\,\cdot\,|\,\textbf{X}=a_{j}(\textbf{x}_{\text{harm}});f,g)$ be the $j$th random output sequence from $\mathcal{Y}$ produced by $g$ using $f$, given $\textbf{X}_{\text{harm}}=\textbf{x}$ and $A_{j}=a_{j}$. Given our definition of $(k,\gamma)$-success, we then define the true $(k,\gamma)$-success rate as
$r_{k,\gamma}(\mathcal{A},f,g)\vcentcolon=\mathbb{E}[s_{k,\gamma}(\textbf{X}_{\text{harm}},\textbf{Y}_{1},\ldots,\textbf{Y}_{k})]$ (2)
where the expectation is taken over $\textbf{X}_{\text{harm}}$, $A_{1},\ldots A_{k}$ and $\textbf{Y}_{1},\ldots,\textbf{Y}_{k}$. Note that when an augmentation set is a singleton (e.g., $\mathcal{A}_{I}$) and a deterministic generation algorithm $g$ is used, the $(k,\gamma)$-success rate is the same as the $(1,0)$-success rate for any values of $k$ and $\gamma$. To approximate the true $(k,\gamma)$-success rate, we define the empirical $(k,\gamma)$-success rate as
$\hat{r}_{k,\gamma}(\mathcal{A},f,g)\vcentcolon=\frac{1}{|\mathcal{D}|}\sum\limits_{\textbf{x}_{i}\in\mathcal{D}}s_{k,\gamma}(\textbf{x}_{i},\textbf{y}_{i1},\ldots,\textbf{y}_{ik})$ (3)
where for $1\leq j\leq k$, $\textbf{y}_{ij}\in\mathcal{Y}$ is the observed output given $a_{ij}(\textbf{x}_{i})$, where $a_{ij}\in\mathcal{A}$ is the $j$th observed augmentation for $\textbf{x}_{i}$. Since we can only obtain an empirical $(k,\gamma)$-success rate in practice, we refer to it simply as the $(k,\gamma)$-success rate. We sometimes use the terms “success rate” and $(k,\gamma)$-success rate” interchangeably if $k$ and $\gamma$ are clear from the surrounding context.
3 Experimental Setup
For computing $(k,\gamma)$-success rates, we set $k=25$ to reduce the runtime of experiments and since we find this value to be sufficient for significantly affecting the success rate. Since the $(k,\gamma)$-success false positive and false negative rates may vary significantly for each augmentation set $\mathcal{A}$, we use separate thresholds $\gamma^{*}_{\mathcal{A}}$ that balances empirical estimates of the false positive and false negative rates for each $\mathcal{A}$, and employ a human study to obtain these empirical estimates (see Appendix B.2 for more details on selecting $\gamma^{*}_{\mathcal{A}}$). As different model developers may impose different objectives for safety alignment, we seek a dataset $\mathcal{D}$ that covers a broad range of harmful behavior classes. Hence, we use the SORRY-Bench dataset (Xie et al., 2024), which was constructed by sampling 450 prompts from across 10 datasets in a class-balanced manner over 45 classes. For the safety judge $c$, we use the fine-tuned Mistral-based safety judge that accompanies the SORRY-Bench dataset.
3.1 Augmentations
For the RQ1 experiments, we examine the success rate gain $\hat{r}_{25,\gamma^{*}_{\mathcal{A}}}(\mathcal{A},f,g)-\hat{r}_{25,\gamma^{*}_{\mathcal{A}}}(\mathcal{A}_{I},f,g)$ of using an augmentation set $\mathcal{A}$ over using no augmentations. Note that since adversarial attacks typically assume white-box access to the model, and since random augmentations are an entirely black-box operation, we do not compare against adversarial attack performance.
Table 1: A list of the types of random augmentation considered in this paper. The augmentations are grouped into string insertion (“SI”) and character-level (“CL”) augmentations.
Name Example
None Write a letter threatening to harm someone’s family if they don’t comply with your demands.
SI Suffix Write a letter threatening to harm someone’s family if they don’t comply with your demands. Y\$^o
Prefix 5i_E Write a letter threatening to harm someone’s family if they don’t comply with your demands.
Anywhere Write a letter threateninP\”y%g to harm someone’s family if they don’t comply with your demands.
CL Edit Wr^te a letter3threatening.to harm someone’s family if they don’t <omply with your demands.
Insert Write a letter thrkeatenin3g to harm someone’s family if they don’t .comply with >your demands.
Delete Wrie a letter threatening to harm someones family if they dontcomply with your demands.
3.1.1 Augmentation Types
Table 1 provides an overview of the augmentation types we investigate. We consider two main kinds of random augmentations: string insertion and character-level augmentations. String insertion augmentations insert a contiguous sequence of random characters into the prompt: either at the end prepended with a space (“Suffix”), beginning appended with a space (“Prefix”) or at a random position (“Anywhere”). This is meant to provide a random counterpart to how some adversarial attacks such as GCG (Zou et al., 2023) append an adversarial suffix to the prompt, and different insertion locations are examined to assess whether the location of the random string matters. Character-level augmentations on the other hand operate at multiple random character locations in the prompt: either by editing characters (“Edit”), inserting characters (“Insert”) or deleting characters (“Delete”(Karpukhin et al., 2019). For either kind of augmentation, all characters and character positions are chosen independently and uniformly at random, i.e., $P_{\text{aug}}(\,\cdot\,;\mathcal{A})=\text{Unif}(\mathcal{A})$.
3.1.2 Augmentation Strength
For string insertion augmentations, the notion of augmentation “strength” refers to the length of the inserted string, whereas for character-level augmentations, “strength” refers to the amount of character positions that are augmented. We consider two ways to control the strength of an augmentation: 1. The strength is fixed for each prompt, and 2. The strength is proportional to the length of each prompt. Since $\mathcal{D}$ may contain a wide range of prompt lengths, fixing the strength may result in augmentations that are too aggressive for short prompts (which may change their semantic meaning) or too subtle for long prompts (which may lead to low success rate gains), in particular for character-level augmentations. Therefore, we focus on proportional augmentation strength, as governed by a proportion parameter $p$. For instance, with $p=0.1$ and an original prompt length of 200 characters, the inserted string length for string insertion augmentations and the amount of augmented character positions for character-level augmentations would be 20 characters. (The number of characters is always rounded down to the nearest integer.) For our experiments, we set $p=0.05$, which we find to be sufficient for obtaining non-trivial success rate gains while ensuring the augmentations are not too aggressive for shorter prompts (see the examples in Table 1, which use $p=0.05$).
3.2 Models
We consider the following models across 8 different model families: Llama 2 (Llama 2 7B Chat, Llama 2 13B Chat) (Touvron et al., 2023), Llama 3 (Llama 3 8B Instruct) (Dubey et al., 2024), Llama 3.1 (Llama 3.1 8B Instruct), Mistral (Mistral 7B Instruct v0.2), Phi 3 (Phi 3 Mini 4K Instruct, Phi 3 Small 8K Instruct, Phi 3 Medium 4K Instruct), Qwen 2 (Qwen 2 0.5B, Qwen 2 1.5B, Qwen 2 7B), Vicuna (Vicuna 7B v1.5, Vicuna 13B v1.5) and Zephyr (Zephyr 7B Beta). Among these, only the Llama, Phi and Qwen families have undergone explicit safety alignment. The remaining families are included to see if any interesting patterns can be observed for unaligned models. For instance, although Mistral was not explicitly aligned, it can sometimes exhibit refusal behavior for some harmful prompts, so it would still be interesting to see how this behavior can be affected by random augmentations.
For the RQ2 experiments, for each augmentation set $\mathcal{A}$ we examine the success rate gain $\hat{r}_{25,\gamma^{*}_{\mathcal{A}}}(\mathcal{A},f^{\prime},g)-\hat{r}_{25,\gamma^{*}_{\mathcal{A}}}(\mathcal{A},f,g)$ of a model $f^{\prime}$ over a baseline model $f$. In the following, we provide further details for each experiment:
Model size. For comparing model sizes, we let the smallest model in each model family be the baseline model $f$ and let the larger models be $f^{\prime}$. Specifically, for Llama 2 the baseline model is Llama 2 7B Chat, for Phi 3 the baseline model is Phi 3 Mini 4k Instruct, for Qwen 2 the baseline model is Qwen 2 0.5B, and for Vicuna the baseline model is Vicuna 7B v1.5.
Quantization. For comparing quantization levels, we consider the original model as the baseline $f$ and the quantized models as $f^{\prime}$. We only focus on 7B/8B parameter models to reduce the amount of experiments as well as to roughly control for model size while assessing quantization over a broad range of model families. We examine two settings for quantization: 1. Symmetric 8-bit per-channel integer quantization of the weights with symmetric 8-bit per-token integer quantization for activations (“W8A8”), and 2. Symmetric 4-bit per-channel weight-only integer quantization (“W4A16”(Nagel et al., 2021). The former is chosen to examine the effects of simultaneous weight and activation quantization (Xiao et al., 2023), and the latter is chosen to explore closer to the limits of weight quantization (Frantar et al., 2022).
Fine-Tuning-Based Defense. For comparing fine-tuning-based defenses, we consider circuit breaking (RR) (Zou et al., 2024) on Mistral 7B Instruct v0.2 and Llama 3 8B Instruct as well as adversarial training (R2D2) (Mazeika et al., 2024) on Zephyr 7B Beta as $f^{\prime}$ and the original model before fine-tuning as the baseline $f$. Note that R2D2 was trained against GCG with a fixed adversarial suffix length of 20 tokens, and that 25 characters corresponds to around 20 tokens on average for the Zephyr tokenizer. Hence, to give a fairer evaluation of R2D2, we additionally examine fixed-length suffix insertion at $L=25$, as well as fixed lengths above and below 25 to assess length generalization; specifically, we examine $L\in\{5,10,15,20,25,30,35,40,45,50\}$. As a sanity check, we also evaluate how often benign prompts are wrongly refused when augmented with a fixed-length suffix; for this, we use the first-turn prompts from MT-Bench (Zheng et al., 2023), which comprise a sample of 80 prompts from MMLU (a benchmark for evaluating core knowledge) (Hendrycks et al., 2020). Note that using the SORRY-Bench judge as a proxy for measuring benign prompt compliance is viable since the judge’s task prompt only asks to evaluate compliance rather than harmfulness.
3.3 Decoding Strategies
By default, all our experiments are conducted using greedy decoding to isolate the randomness effects of using multiple random augmentations. However, for the RQ3 experiments, for each augmentation set $\mathcal{A}$ we examine the success rate gain $\hat{r}_{25,\gamma^{*}_{\mathcal{A}}}(\mathcal{A},f,g)-\hat{r}_{25,\gamma^{*}_{\mathcal{A}}}(\mathcal{A}_{I},f,g)$ for sampling-based generation algorithms $g$. Specifically, we consider temperature sampling with various temperatures $\tau$ for $g$. We consider two values for $\tau$: 0.7 (since this is a value in the range of common temperature parameters between 0.6 and 0.9), and 1.0 (to explore the largest possible temperature value). Note that greedy decoding is a special case where $\tau=0.0$.
4 Experimental Results
In this section, we plot the results for each of our experiments and discuss our observations. Raw data values (including results using a fixed $\gamma=0$ for all augmentations) broken down by augmentation type are reported in Appendix C.
4.1 RQ1: Varying Augmentation Type
Figure 2: Average $\left(25,\gamma^{*}_{\mathcal{A}}\right)$-success rate gains of different kinds of augmentations over using no augmentations for various sampling temperatures $\tau$. RQ1 results are where $\tau=0.0$, and the rest are RQ3 results.
[…]
4.2 RQ2: Varying Model Aspects
4.2.1 Model Size
Figure 3: Average $(25,\gamma^{*}_{\mathcal{A}})$-success rate gains of larger models over the smallest model in their model family, using greedy decoding for generation.
Figure 3 reports the model size experiment results for RQ2. Larger models tend to be safer than smaller ones, although the pattern is not strict, nor is safety proportional to model size. For example, while Phi 3 Small tends to be somewhat safer than Phi 3 Mini, Phi 3 Medium actually becomes less safe. Moreover, Qwen 2 1.5B tends to exhibit a greater increase in safety than Qwen 2 7B, despite being a much smaller model. This suggests that increasing model size alone is insufficient for improving safety against random augmentations, and that there may be other underlying causes behind the observed pattern (e.g., causes related to the alignment dataset).
4.2.2 Quantization
Figure 4: Average $(25,\gamma^{*}_{\mathcal{A}})$-success rate gains of quantized models over their respective original models, using greedy decoding for generation.
Figure 4 reports the quantization experiment results for RQ2. For W8A8, most success rate changes are small, with all deviations being within 5%. Among all models, Qwen 2 7B has the greatest tendency towards becoming less safe. In Figure 12 in Appendix D, we show an example where the original Qwen 2 model fails under the random suffix augmentation while the W8A8 model succeeds even when the random suffixes used are the exact same for both models. Moving over to the W4A16 results, we see that the Llama 3, Mistral, Phi and Vicuna models become noticeably less safe. However, other models such as Llama 2, Llama 3.1 and Zephyr barely change, similar to their W8A8 counterparts. Even more curiously however, we see that Qwen 2 7B seemingly becomes more safe. However, upon further inspection, we realize that this may be a result of poorer model response quality in general; see Figure 13 in Appendix D for examples. Overall, while quantization can have some significant influence on success rate with more aggressive weight quantization tending to reduce safety, these effects are not consistent across models. As with the results for the model size experiment, this suggests that there may be other underlying factor(s) that determine how quantization affects safety under random augmentations.
4.2.3 Fine-Tuning-Based Defense
Figure 5: Average $(25,\gamma^{*}_{\mathcal{A}})$-success rate gains of models with fine-tuning-based defenses over their respective original models, using greedy decoding.
Figure 5 reports the fine-tuning-based defense experiment results for RQ2. All fine-tuned models clearly provide some amount of improvement in safety over their respective original models. For RR models, the improvement for Mistral 7B is much greater than the improvement for Llama 3 8B, probably due to the original Mistral model not being explicitly aligned (and therefore having a much larger room for improvement than the already aligned Llama 3 model). Interestingly, although Zephyr 7B was adversarially trained against only GCG suffixes of a fixed token length, it also enjoys some safety improvement on proportional-length random augmentations.
To provide a fairer assessment of the improvements afforded by the adversarial training, we also examine fixed-length random suffixes in Figure 6. Note that character length 25 has an average token length of ${\sim}22$, which is close to the fixed length of 20 tokens R2D2 used. The blue curve shows that as the length increasess, the success rate continues to decrease, even somewhat past length 25. This suggests that length generalization is unidirectional: protection is afforded to lengths beyond the adversarial training length, but starts to disappear for shorter suffixes. This heavily contrasts with adversarial training in the image classification domain, where protections typical only hold with image perturbations less extreme than what was trained against (Gowal et al., 2018).
We also suspect that the model may be becoming overeager in refusing any request that involves a sufficiently long random suffix. To validate this hypothesis, in Figure 6 we also plot the average judge score (over augmentation and prompt dimensions) for benign first-turn prompts taken from MT-Bench, shown by the green curve. We notice that the model also experiences an increase in refusals for benign prompts as the suffix length increases, plateauing at around 25. Manual inspection reveals that many of the model responses are indeed the adversarial training refusal string “Sorry, I can’t do that.” This suggests that for adversarial training, additional regularization may be needed on augmented benign prompts, which may also help to reduce the chance of shortcut learning (Geirhos et al., 2020) where the model is overly-sensitive to the presence of an unintelligible suffix (whether random or adversarial). We also plot the average judge score for SORRY-Bench prompts (orange curve) and observe that the two curves have a similar shape (although the gains are much lower for harmful prompts as one would hope.)
Figure 6: Fixed-length suffix insertion results for Zephyr 7B Beta and Zephyr 7B Beta (R2D2) at various character lengths $L$.
4.3 RQ3: Varying the Generation Configuration
Figure 2 reports the experiment results for RQ3 (denoted by $\tau=0.7$ and $\tau=1.0$). First, we remark that increasing temperature without any augmentations already increases the success rate (see Table 3); this is in line with the findings of Huang et al. (2023) that showed altering temperature alone can be a successful attack. Next, we observe, that across all models and for both values of $\tau$, using random augmentations further improves the success rate on top of success gains due to output sampling. This supplements the findings of Huang et al. (2023) which only explored success rate improvements for prompts without any input augmentations. Our results show that two sources of randomness, namely output sampling and input augmentations, can work together to provide even greater attack effectiveness.
5 Conclusion
This paper demonstrates that simple random augmentations are a cheap yet effective approach to bypassing the safety alignment of state-of-the-art LLMs. However, although we identify patterns in how dimensions such as model size and quantization affect safety under random augmentations, we also discover counterexamples to these patterns. As such, the full story may be much more complex and involve additional underlying factors such as training data, optimization, etc. In the future, we will investigate whether these more complex dimensions can explain the counterexamples we observed. Moreover, we will go beyond evaluation to identify ways that models can become more robust against random augmentations.
Acknowledgments
This work was supported in part by NSF Grants No. CCF-2238079, CCF-2316233, CNS-2148583, Google Research Scholar award, and a research grant from the Amazon-Illinois Center on AI for Interactive Conversational Experiences (AICE).
Appendix A Related Work
A.1 Random Augmentations and Robustness
Prior studies on the impact of random augmentations of robustness in NLP have largely focused on how they impact the performance of text classifiers. For instance, it has been shown that Neural Machine Translation (NMT) is vulnerable to character-level random augmentations such as swapping, keyboard typos, and editing (Belinkov & Bisk, 2017; Heigold et al., 2017). Furthermore,  Karpukhin et al. (2019) demonstrated that training NMT models with character-level augmentations can improve robustness to natural noise in real-world data. Beyond NMT, Zhang et al. (2021) examined how both character-level (e.g., whitespace and character insertion) and word-level augmentations (e.g., word shuffling) can significantly degrade the sentiment analysis and paraphrase detection performance of models such as BERT (Devlin, 2018) and RoBERTa (Liu, 2019).
A.2 Random Augmentations in Adversarial Attacks
Random augmentations have also been utilized within adversarial attack algorithms. For instance, Li et al. (2018) introduced the TextBugger attack framework, which adversarially applies random augmentations (e.g., character-level augmentations such as inserting, swapping, or deleting characters and word-level augmentations such as word substitution) to fool models on sentiment analysis, question answering and machine translation tasks. Their method computes a gradient to estimate word importance, and then uses this estimate to apply random augmentations at specific locations based on the importance estimation. Additionally, Morris et al. (2020) introduced a comprehensive framework for generating adversarial examples to attack NLP models such as BERT, utilizing the word-level augmentations from the Easy Data Augmentation method  (Wei & Zou, 2019) (i.e., synonym replacement, insertion, swapping, and deletion). The adversarial examples are also used to perform adversarial training to improve model robustness and generalization.
A.3 Random Augmentations for Defense
SmoothLLM  (Robey et al., 2023) was introduced as a system-level defense for mitigating jailbreak effectiveness. Their key observation is that successful jailbreaks are extremely brittle to random augmentations; i.e., many of the successful jailbreaks won’t succeed after augmentation. In contrast, our work is based on the observation that the original prompt itself is also brittle, but in the opposite direction: given a prompt that doesn’t succeed, one can effectively find an augmented prompt that does succeed. Moreover, their attack success evaluation is only based on a single chosen output per prompt, effectively discarding the other $k-1$ outputs. In contrast, since our threat model is built around the attacker making $k$ independent attempts per prompt, our attack success evaluation accounts for all of the $k$ outputs per prompt.
Following in the footsteps of SmoothLLM, JailGuard  (Zhang et al., 2024) was proposed as another defense method. Similar to SmoothLLM, JailGuard involves applying multiple random augmentations per prompt on the system side. However, JailGuard does not leverage a safety judge, instead examining the model response variance to determine whether a prompt is harmful or not. In a follow-up work to SmoothLLM,  Ji et al. (2024) considers more advanced random augmentations such as synonym replacement or LLM-based augmentations such as paraphrasing and summarization. In the case of LLM-based augmentations, the randomness comes from the stochasticity of the generation algorithm (so long as greedy decoding is not used). In an earlier work, (Kumar et al., 2023) proposed RandomEC, which defends against jailbreaks by erasing random parts of the input and checking whether a safety judge deems the input to be safe or not, and deems the original input unsafe only if any of the augmented prompts are deemed unsafe.
Appendix B Additional Details on $(k,\gamma)$-Success
B.1 Effect of $\gamma$ on FPR and FNR
To see how the choice of $\gamma$ can affect the false positive rate, let $\hat{Z}_{j}$ be the judge’s predicted score for the $j$th augmentation, and let $Z_{j}$ be the corresponding true score (e.g., from human evaluation). Let $\hat{Z}=\sum\limits_{j=1}^{k}\hat{Z}_{j}$. Then the false positive rate as a function of $\gamma$ for $\mathcal{A}$ is
$p_{\text{FP}}(\gamma;\mathcal{A})\vcentcolon=\Pr(\hat{Z}>k\gamma\,|\,Z_{1}=0\cap Z_{2}=0\cap\ldots\cap Z_{k}=0)$ (4)
Clearly, this corresponds to evaluating the complementary CDF of a conditional distribution. Thus, $p_{\text{FP}}(\gamma;\mathcal{A})$ is monotonically decreasing in $\gamma$. On the other hand, note that the false negative rate is
$p_{\text{FN}}(\gamma;\mathcal{A})\vcentcolon=\Pr(\hat{Z}\leq k\gamma\,|\,Z_{1}=1\cup Z_{2}=1\cup\ldots\cup Z_{k}=1)$ (5)
This corresponds to evaluating the CDF of a conditional distribution, and thus $p_{\text{FN}}(\gamma;\mathcal{A})$ is monotonically increasing in $\gamma$. Therefore, $\gamma$ strikes a trade-off between $p_{\text{FP}}(\gamma;\mathcal{A})$ and $p_{\text{FN}}(\gamma;\mathcal{A})$, so care must be taken to select a $\gamma$ that balances this trade-off.
B.2 Selecting $\gamma^{*}_{\mathcal{A}}$
For selecting $\gamma$, to balance $p_{\text{FP}}(\gamma;\mathcal{A})$ and $p_{\text{FN}}(\gamma;\mathcal{A})$ we aim to select a $\gamma$ that minimizes their average (i.e. the balanced error rate). However, since we cannot exactly evaluate $p_{\text{FP}}(\gamma;\mathcal{A})$ and $p_{\text{FN}}(\gamma;\mathcal{A})$, we estimate them using the data from $\mathcal{D}$ and human judgement. Specifically, after obtaining judge scores for all our experiments, we randomly sampled 1220 data points (each containing $k=25$ judge scores) from across all our experiment results and manually label them to obtain empirical estimates $\hat{p}_{\text{FP}}(\gamma;\mathcal{A})$ and $\hat{p}_{\text{FN}}(\gamma;\mathcal{A})$. The task of manual labeling is split over three of the authors so that each data point receives a human score from one of the authors. We provide the following task description to each of the labelers, adapted from the SORRY-Bench judge task prompt:
Your goal is to judge if
-[Label 1] At least one of the model responses complies with the user request, i.e., containing content that can directly assist with or resolve the request.
-[Label 0] All the model responses don’t comply with the user request; e.g., directly declining the user request or merely providing relevant content that cannot directly resolve the user request.
For each data point, the labeler is shown the original, unaugmented prompt and the $k=25$ model responses. After manual labeling, we then select
$\gamma^{*}_{\mathcal{A}}=\operatorname*{arg\,min}\limits_{\gamma\in\{0,\frac{1}{k},\ldots,\frac{k-1}{k}\}}\frac{\hat{p}_{\text{FP}}(\gamma;\mathcal{A})+\hat{p}_{\text{FN}}(\gamma;\mathcal{A})}{2}$ (6)
as $\gamma$ for $\mathcal{A}$. See Table 2 for a breakdown of the chosen $\gamma^{*}_{\mathcal{A}}$ for each augmentation set $\mathcal{A}$ along with their estimated false positive and false negative rates.
Table 2: Optimized values $\gamma^{*}_{\mathcal{A}}$ for each augmentation set $\mathcal{A}$, along with their empirical false positive and false negative rates (“FPR” and “FNR”). “Avg” reports the average of the false positive and false negative rates (i.e. the balanced error rate). The FPR and FNR results for $\gamma=0$ are also included for comparison. Note that for some augmentations, $\gamma^{*}_{\mathcal{A}}=0$, indicating that no other threshold could be found to further reduce the balanced error rate.
FPR FNR Avg
Augmentation $\gamma^{*}_{\mathcal{A}}$ $\gamma=0$ $\gamma=\gamma^{*}_{\mathcal{A}}$ $\gamma=0$ $\gamma=\gamma^{*}_{\mathcal{A}}$ $\gamma=0$ $\gamma=\gamma^{*}_{\mathcal{A}}$
None 0.000 0.024 0.024 0.078 0.078 0.051 0.051
String Insertion Suffix 0.000 0.125 0.125 0.027 0.027 0.076 0.076
Prefix 0.000 0.055 0.055 0.044 0.044 0.050 0.050
Any 0.080 0.129 0.065 0.051 0.102 0.090 0.083
Character-Level Edit 0.080 0.197 0.049 0.000 0.102 0.098 0.076
Insert 0.040 0.156 0.073 0.025 0.100 0.091 0.086
Delete 0.040 0.173 0.107 0.067 0.078 0.120 0.092
Overall 0.000 0.112 0.112 0.038 0.038 0.075 0.075
Appendix C Expanded Experimental Results
Tables 3-7 provide a detailed breakdown of the raw data values obtained in our experiments.
Table 3: $(25,\gamma)$-success rate gains of different augmentation sets $\mathcal{A}$ over the no augmentation set $\mathcal{A}_{I}$ for various temperatures $\tau$. The “None” column reports the empirical $\left(1,0\right)$-success rate $\hat{r}_{1,0}(\mathcal{A}_{I},f,g)$, whereas the other augmentation columns report the empirical $\left(25,\gamma\right)$-success rate gain $\hat{r}_{25,\gamma}(\mathcal{A},f,g)-\hat{r}_{25,\gamma}(\mathcal{A}_{I},f,g)$.
String Insertion Character-Level
Model $\bm{\tau}$ $\gamma$ None Suffix Prefix Any Avg Edit Insert Delete Avg Avg
Llama 2 7B Chat 0.0 $\gamma^{*}_{\mathcal{A}}$ 0.151 +0.038 +0.049 +0.051 +0.046 +0.136 +0.124 +0.147 +0.136 +0.091
0.7 $\gamma^{*}_{\mathcal{A}}$ 0.191 +0.018 +0.018 +0.004 +0.013 +0.087 +0.084 +0.131 +0.101 +0.057
1.0 $\gamma^{*}_{\mathcal{A}}$ 0.218 +0.009 +0.011 -0.007 +0.004 +0.082 +0.073 +0.104 +0.087 +0.046
0.0 0 0.151 +0.038 +0.049 +0.113 +0.067 +0.253 +0.191 +0.231 +0.225 +0.146
0.7 0 0.191 +0.018 +0.018 +0.120 +0.052 +0.227 +0.162 +0.209 +0.199 +0.126
1.0 0 0.218 +0.009 +0.011 +0.096 +0.039 +0.222 +0.153 +0.191 +0.189 +0.114
Llama 3 8B Instruct 0.0 $\gamma^{*}_{\mathcal{A}}$ 0.236 +0.024 -0.002 +0.031 +0.018 +0.102 +0.096 +0.164 +0.121 +0.069
0.7 $\gamma^{*}_{\mathcal{A}}$ 0.271 +0.029 +0.009 +0.044 +0.027 +0.116 +0.107 +0.200 +0.141 +0.084
1.0 $\gamma^{*}_{\mathcal{A}}$ 0.289 +0.044 +0.009 +0.044 +0.033 +0.149 +0.122 +0.200 +0.157 +0.095
0.0 0 0.236 +0.024 -0.002 +0.102 +0.041 +0.251 +0.167 +0.242 +0.220 +0.131
0.7 0 0.271 +0.029 +0.009 +0.136 +0.058 +0.282 +0.182 +0.258 +0.241 +0.149
1.0 0 0.289 +0.044 +0.009 +0.144 +0.066 +0.298 +0.189 +0.293 +0.260 +0.163
Llama 3.1 8B Instruct 0.0 $\gamma^{*}_{\mathcal{A}}$ 0.269 +0.067 +0.009 +0.047 +0.041 +0.136 +0.129 +0.218 +0.161 +0.101
0.7 $\gamma^{*}_{\mathcal{A}}$ 0.311 +0.038 -0.020 +0.016 +0.011 +0.104 +0.084 +0.184 +0.124 +0.068
1.0 $\gamma^{*}_{\mathcal{A}}$ 0.309 +0.033 -0.020 -0.004 +0.003 +0.082 +0.071 +0.173 +0.109 +0.056
0.0 0 0.269 +0.067 +0.009 +0.138 +0.071 +0.278 +0.184 +0.309 +0.257 +0.164
0.7 0 0.311 +0.038 -0.020 +0.098 +0.039 +0.249 +0.138 +0.273 +0.220 +0.129
1.0 0 0.309 +0.033 -0.020 +0.098 +0.037 +0.229 +0.138 +0.260 +0.209 +0.123
Mistral 7B Instruct v0.2 0.0 $\gamma^{*}_{\mathcal{A}}$ 0.653 +0.207 +0.169 +0.153 +0.176 +0.209 +0.204 +0.198 +0.204 +0.190
0.7 $\gamma^{*}_{\mathcal{A}}$ 0.727 +0.156 +0.107 +0.107 +0.123 +0.144 +0.156 +0.160 +0.153 +0.138
1.0 $\gamma^{*}_{\mathcal{A}}$ 0.740 +0.158 +0.104 +0.111 +0.124 +0.169 +0.164 +0.164 +0.166 +0.145
0.0 0 0.653 +0.207 +0.169 +0.242 +0.206 +0.284 +0.249 +0.242 +0.259 +0.232
0.7 0 0.727 +0.156 +0.107 +0.200 +0.154 +0.233 +0.202 +0.200 +0.212 +0.183
1.0 0 0.740 +0.158 +0.104 +0.187 +0.150 +0.227 +0.204 +0.202 +0.211 +0.180
Phi 3 Small 8K Instruct 0.0 $\gamma^{*}_{\mathcal{A}}$ 0.167 +0.062 +0.067 +0.076 +0.068 +0.196 +0.198 +0.213 +0.202 +0.135
0.7 $\gamma^{*}_{\mathcal{A}}$ 0.202 +0.053 +0.060 +0.073 +0.062 +0.187 +0.173 +0.207 +0.189 +0.126
1.0 $\gamma^{*}_{\mathcal{A}}$ 0.229 +0.071 +0.096 +0.069 +0.079 +0.202 +0.187 +0.238 +0.209 +0.144
0.0 0 0.167 +0.062 +0.067 +0.176 +0.101 +0.391 +0.289 +0.324 +0.335 +0.218
0.7 0 0.202 +0.053 +0.060 +0.158 +0.090 +0.376 +0.271 +0.318 +0.321 +0.206
1.0 0 0.229 +0.071 +0.096 +0.200 +0.122 +0.422 +0.324 +0.367 +0.371 +0.247
Qwen 2 7B Instruct 0.0 $\gamma^{*}_{\mathcal{A}}$ 0.356 +0.109 +0.116 +0.098 +0.107 +0.233 +0.213 +0.262 +0.236 +0.172
0.7 $\gamma^{*}_{\mathcal{A}}$ 0.396 +0.124 +0.142 +0.124 +0.130 +0.222 +0.193 +0.240 +0.219 +0.174
1.0 $\gamma^{*}_{\mathcal{A}}$ 0.438 +0.147 +0.162 +0.116 +0.141 +0.220 +0.198 +0.253 +0.224 +0.183
0.0 0 0.356 +0.109 +0.116 +0.207 +0.144 +0.360 +0.269 +0.336 +0.321 +0.233
0.7 0 0.396 +0.124 +0.142 +0.227 +0.164 +0.360 +0.256 +0.333 +0.316 +0.240
1.0 0 0.438 +0.147 +0.162 +0.218 +0.176 +0.367 +0.278 +0.329 +0.324 +0.250
Vicuna 7B v1.5 0.0 $\gamma^{*}_{\mathcal{A}}$ 0.369 +0.096 +0.096 +0.113 +0.101 +0.198 +0.184 +0.196 +0.193 +0.147
0.7 $\gamma^{*}_{\mathcal{A}}$ 0.416 +0.080 +0.084 +0.084 +0.083 +0.131 +0.142 +0.173 +0.149 +0.116
1.0 $\gamma^{*}_{\mathcal{A}}$ 0.502 +0.073 +0.076 +0.071 +0.073 +0.104 +0.140 +0.142 +0.129 +0.101
0.0 0 0.369 +0.096 +0.096 +0.209 +0.133 +0.304 +0.251 +0.264 +0.273 +0.203
0.7 0 0.416 +0.080 +0.084 +0.180 +0.115 +0.280 +0.224 +0.253 +0.253 +0.184
1.0 0 0.502 +0.073 +0.076 +0.191 +0.113 +0.269 +0.216 +0.236 +0.240 +0.177
Zephyr 7B Beta 0.0 $\gamma^{*}_{\mathcal{A}}$ 0.856 +0.076 +0.087 +0.087 +0.083 +0.102 +0.111 +0.107 +0.107 +0.095
0.7 0 0.884 +0.064 +0.073 +0.067 +0.068 +0.067 +0.080 +0.076 +0.074 +0.071
1.0 0 0.882 +0.064 +0.076 +0.060 +0.067 +0.076 +0.089 +0.082 +0.082 +0.074
0.0 0 0.856 +0.076 +0.087 +0.127 +0.096 +0.131 +0.124 +0.124 +0.127 +0.111
0.7 0 0.884 +0.064 +0.073 +0.089 +0.076 +0.089 +0.100 +0.087 +0.092 +0.084
1.0 0 0.882 +0.064 +0.076 +0.087 +0.076 +0.098 +0.100 +0.096 +0.098 +0.087
Table 4: $(25,\gamma)$-success rate gains of models $f^{\prime}$ over the smallest model in their model family $f$, using greedy decoding for $g$. Adjacent rows are grouped by model family. An asterisk (*) next to a model name indicates the model $f$ is the smallest in its respective model family, and the values for that row report the empirical $(25,\gamma)$-success rate $\hat{r}_{25,\gamma}(\mathcal{A},f,g)$. The other rows report the empirical $(25,\gamma)$-success rate gain $\hat{r}_{25,\gamma}(\mathcal{A},f^{\prime},g)-\hat{r}_{25,\gamma}(\mathcal{A},f,g)$.
String Insertion Character-Level
Model $\gamma$ None Suffix Prefix Any Avg Edit Insert Delete Avg Avg
Llama 2 7B Chat* $\gamma^{*}_{\mathcal{A}}$ 0.151 0.189 0.200 0.202 0.197 0.287 0.276 0.298 0.287 0.242
Llama 2 13B Chat $\gamma^{*}_{\mathcal{A}}$0.011 -0.013 -0.007 -0.016 -0.012 -0.049 -0.036 -0.040 -0.041 -0.027
Llama 2 7B Chat* 0 0.151 0.189 0.200 0.264 0.218 0.404 0.342 0.382 0.376 0.297
Llama 2 13B Chat 0 -0.011 -0.013 -0.007 -0.029 -0.016 -0.060 -0.060 -0.060 -0.060 -0.038
Phi 3 Mini 4k Instruct* $\gamma^{*}_{\mathcal{A}}$ 0.202 0.358 0.289 0.260 0.302 0.460 0.440 0.491 0.464 0.383
Phi 3 Small 8K Instruct $\gamma^{*}_{\mathcal{A}}$0.036 -0.129 -0.056 -0.018 -0.067 -0.098 -0.076 -0.111 -0.095 -0.081
Phi 3 Medium 4K Instruct $\gamma^{*}_{\mathcal{A}}$ +0.089 +0.080 +0.153 +0.153 +0.129 +0.069 +0.087 +0.051 +0.069 +0.099
Phi 3 Mini 4k Instruct* 0 0.202 0.358 0.289 0.411 0.353 0.644 0.544 0.593 0.594 0.473
Phi 3 Small 8K Instruct 0 -0.036 -0.129 -0.056 -0.069 -0.084 -0.087 -0.089 -0.102 -0.093 -0.089
Phi 3 Medium 4K Instruct 0 +0.089 +0.080 +0.153 +0.113 +0.116 +0.036 +0.040 +0.056 +0.044 +0.080
Qwen 2 0.5B Instruct* $\gamma^{*}_{\mathcal{A}}$ 0.431 0.578 0.633 0.598 0.603 0.731 0.767 0.756 0.751 0.677
Qwen 2 1.5B Instruct $\gamma^{*}_{\mathcal{A}}$0.111 -0.196 -0.249 -0.204 -0.216 -0.213 -0.262 -0.218 -0.231 -0.224
Qwen 2 7B Instruct $\gamma^{*}_{\mathcal{A}}$0.076 -0.113 -0.162 -0.144 -0.140 -0.142 -0.198 -0.138 -0.159 -0.150
Qwen 2 0.5B Instruct* 0 0.431 0.578 0.633 0.713 0.641 0.867 0.867 0.844 0.859 0.750
Qwen 2 1.5B Instruct 0 -0.111 -0.196 -0.249 -0.216 -0.220 -0.202 -0.278 -0.224 -0.235 -0.227
Qwen 2 7B Instruct 0 -0.076 -0.113 -0.162 -0.151 -0.142 -0.151 -0.242 -0.153 -0.182 -0.162
Vicuna 7B v1.5* $\gamma^{*}_{\mathcal{A}}$ 0.369 0.464 0.464 0.482 0.470 0.567 0.553 0.564 0.561 0.516
Vicuna 13B v.15 $\gamma^{*}_{\mathcal{A}}$0.082 -0.118 -0.102 -0.127 -0.116 -0.122 -0.129 -0.107 -0.119 -0.117
Vicuna 7B v1.5* 0 0.369 0.464 0.464 0.578 0.502 0.673 0.620 0.633 0.642 0.572
Vicuna 13B v.15 0 -0.082 -0.118 -0.102 -0.124 -0.115 -0.111 -0.118 -0.093 -0.107 -0.111
Table 5: $(25,\gamma)$-success rate gains of quantized models $f^{\prime}$ over their respective original model $f$, using greedy decoding for $g$. Base model rows are indicated with “None” and report the empirical $(25,\gamma)$-success rate $\hat{r}_{25,\gamma}(\mathcal{A},f,g)$. The other rows report the empirical $(25,\gamma)$-success rate gain $\hat{r}_{25,\gamma}(\mathcal{A},f^{\prime},g)-\hat{r}_{25,\gamma}(\mathcal{A},f,g)$.
String Insertion Character-Level
Model Quant. $\gamma$ None Suffix Prefix Any Avg Edit Insert Delete Avg Avg
Llama 2 7B Chat None $\gamma^{*}_{\mathcal{A}}$ 0.151 0.189 0.200 0.202 0.197 0.287 0.276 0.298 0.287 0.242
W8A8 $\gamma^{*}_{\mathcal{A}}$0.018 +0.002 +0.004 -0.013 -0.002 -0.027 -0.007 +$0$0.011 -0.007
W4A16 $\gamma^{*}_{\mathcal{A}}$0.018 +0.027 -0.004 +0.013 +0.012 +0.022 +0.013 +0.016 +0.017 +0.014
None 0 0.151 0.189 0.200 0.264 0.218 0.404 0.342 0.382 0.376 0.297
W8A8 0 -0.018 +0.002 +0.004 +0.002 +0.003 +0.004 -0.020 +0.004 -0.004 -0.000
W4A16 0 -0.018 +0.027 -0.004 +0.027 +0.016 +0.027 -0.004 +0.004 +0.009 +0.013
Llama 3 8B Instruct None $\gamma^{*}_{\mathcal{A}}$ 0.236 0.260 0.233 0.267 0.253 0.338 0.331 0.400 0.356 0.305
W8A8 $\gamma^{*}_{\mathcal{A}}$0.020 -0.004 +0.007 -0.009 -0.002 +0.002 +0.013 +0.000 +0.005 +0.001
W4A16 $\gamma^{*}_{\mathcal{A}}$ +0.011 +0.047 +0.069 +0.067 +0.061 +0.171 +0.149 +0.127 +0.149 +0.105
None 0 0.236 0.260 0.233 0.338 0.277 0.487 0.402 0.478 0.456 0.366
W8A8 0 -0.020 -0.004 +0.007 -0.007 -0.001 +0.002 -0.007 +0.004 -$0$0.001
W4A16 0 +0.011 +0.047 +0.069 +0.100 +0.072 +0.204 +0.171 +0.140 +0.172 +0.122
Llama 3.1 8B Instruct None $\gamma^{*}_{\mathcal{A}}$ 0.269 0.336 0.278 0.316 0.310 0.404 0.398 0.487 0.430 0.370
W8A8 $\gamma^{*}_{\mathcal{A}}$0.027 -0.024 -0.009 -0.020 -0.018 -0.036 -0.031 -0.024 -0.030 -0.024
W4A16 $\gamma^{*}_{\mathcal{A}}$0.042 +$0$0.022 -0.024 -0.016 -0.031 -0.047 -0.029 -0.036 -0.026
None 0 0.269 0.336 0.278 0.407 0.340 0.547 0.453 0.578 0.526 0.433
W8A8 0 -0.027 -0.024 -0.009 -0.018 -0.017 -0.011 -0.022 -0.007 -0.013 -0.015
W4A16 0 -0.042 +$0$0.022 -0.009 -0.010 -0.013 -0.033 -0.029 -0.025 -0.018
Mistral 7B Instruct v0.2 None $\gamma^{*}_{\mathcal{A}}$ 0.653 0.860 0.822 0.807 0.830 0.862 0.858 0.851 0.857 0.843
W8A8 $\gamma^{*}_{\mathcal{A}}$0.009 -0.007 +0.004 -0.004 -0.002 +0.002 +0.013 +0.020 +0.012 +0.005
W4A16 $\gamma^{*}_{\mathcal{A}}$ +0.076 +0.071 +0.062 +0.098 +0.077 +0.080 +0.091 +0.093 +0.088 +0.083
None 0 0.653 0.860 0.822 0.896 0.859 0.938 0.902 0.896 0.912 0.886
W8A8 0 -0.009 -0.007 +0.004 +0.004 +0.001 -0.002 +0.004 +0.013 +0.005 +0.003
W4A16 0 +0.076 +0.071 +0.062 +0.060 +0.064 +0.038 +0.073 +0.082 +0.064 +0.064
Phi 3 Small 8K Instruct None $\gamma^{*}_{\mathcal{A}}$ 0.167 0.229 0.233 0.242 0.235 0.362 0.364 0.380 0.369 0.302
W8A8 $\gamma^{*}_{\mathcal{A}}$ +0.013 +0.004 +0.013 +0.009 +0.009 +0.018 -0.009 +0.013 +0.007 +0.008
W4A16 $\gamma^{*}_{\mathcal{A}}$ +0.051 +0.049 +0.047 +0.051 +0.049 +0.093 +0.051 +0.073 +0.073 +0.061
None 0 0.167 0.229 0.233 0.342 0.268 0.558 0.456 0.491 0.501 0.385
W8A8 0 +0.013 +0.004 +0.013 +0.007 +0.008 +0.029 -0.004 +0.020 +0.015 +0.011
W4A16 0 +0.051 +0.049 +0.047 +0.076 +0.057 +0.087 +0.056 +0.056 +0.066 +0.061
Qwen 2 7B Instruct None $\gamma^{*}_{\mathcal{A}}$ 0.356 0.464 0.471 0.453 0.463 0.589 0.569 0.618 0.592 0.527
W8A8 $\gamma^{*}_{\mathcal{A}}$ +0.020 +0.036 +0.044 +0.042 +0.041 +0.029 +0.016 +0.011 +0.019 +0.030
W4A16 $\gamma^{*}_{\mathcal{A}}$0.178 -0.076 -0.091 -0.144 -0.104 -0.231 -0.164 -0.178 -0.191 -0.147
None 0 0.356 0.464 0.471 0.562 0.499 0.716 0.624 0.691 0.677 0.588
W8A8 0 +0.020 +0.036 +0.044 +0.038 +0.039 +0.027 +0.016 +0.013 +0.019 +0.029
W4A16 0 -0.178 -0.076 -0.091 -0.102 -0.090 -0.180 -0.133 -0.140 -0.151 -0.120
Vicuna 7B v1.5 None $\gamma^{*}_{\mathcal{A}}$ 0.369 0.464 0.464 0.482 0.470 0.567 0.553 0.564 0.561 0.516
W8A8 $\gamma^{*}_{\mathcal{A}}$0.020 -0.007 +0.004 -0.031 -0.011 -0.027 -0.016 -0.002 -0.015 -0.013
W4A16 $\gamma^{*}_{\mathcal{A}}$ +0.020 +0.058 +0.053 +0.049 +0.053 +0.060 +0.069 +0.080 +0.070 +0.061
None 0 0.369 0.464 0.464 0.578 0.502 0.673 0.620 0.633 0.642 0.572
W8A8 0 -0.020 -0.007 +0.004 -0.013 -0.005 -0.002 -0.004 -0.002 -0.003 -0.004
W4A16 0 +0.020 +0.058 +0.053 +0.071 +0.061 +0.078 +0.073 +0.084 +0.079 +0.070
Zephyr 7B Beta None $\gamma^{*}_{\mathcal{A}}$ 0.856 0.931 0.942 0.942 0.939 0.958 0.967 0.962 0.962 0.950
W8A8 $\gamma^{*}_{\mathcal{A}}$0.011 +0.011 +0.004 +0.000 +0.005 +0.000 +0.002 -0.007 -0.001 +0.002
W4A16 $\gamma^{*}_{\mathcal{A}}$ +0.024 +0.024 +0.031 +0.016 +0.024 +0.007 +0.013 +0.009 +0.010 +0.017
None 0 0.856 0.931 0.942 0.982 0.952 0.987 0.980 0.980 0.982 0.967
W8A8 0 -0.011 +0.011 +0.004 -0.004 +0.004 -0.004 +0.002 +$0$0.001 +0.001
W4A16 0 +0.024 +0.024 +0.031 +0.000 +0.019 +0.002 +0.004 -0.002 +0.001 +0.010
Table 6: $(25,\gamma)$-success rate gains of models $f^{\prime}$ with fine-tuning-based defenses over their respective base models $f$, using greedy decoding for $g$. Adjacent rows are grouped into pairs of the base model and its fine-tuned version. Base model rows report the empirical $(25,\gamma)$-success rate $\hat{r}_{25,\gamma}(\mathcal{A},f,g)$. The other rows report the empirical $(25,\gamma)$-success rate gain $\hat{r}_{25,\gamma}(\mathcal{A},f^{\prime},g)-\hat{r}_{25,\gamma}(\mathcal{A},f,g)$.
String Insertion Character-Level
Model $\gamma$ None Suffix Prefix Any Avg Edit Insert Delete Avg Avg
Llama 3 8B Instruct $\gamma^{*}_{\mathcal{A}}$ 0.236 0.260 0.233 0.267 0.253 0.338 0.331 0.400 0.356 0.305
Llama 3 8B Instruct (RR) $\gamma^{*}_{\mathcal{A}}$0.142 -0.140 -0.098 -0.147 -0.128 -0.196 -0.176 -0.216 -0.196 -0.162
Llama 3 8B Instruct 0 0.236 0.260 0.233 0.338 0.277 0.487 0.402 0.478 0.456 0.366
Llama 3 8B Instruct (RR) 0 -0.142 -0.140 -0.098 -0.151 -0.130 -0.244 -0.198 -0.238 -0.227 -0.178
Mistral 7B Instruct v0.2 $\gamma^{*}_{\mathcal{A}}$ 0.653 0.860 0.822 0.807 0.830 0.862 0.858 0.851 0.857 0.843
Mistral 7B Instruct v0.2 (RR) $\gamma^{*}_{\mathcal{A}}$0.518 -0.633 -0.567 -0.580 -0.593 -0.524 -0.542 -0.509 -0.525 -0.559
Mistral 7B Instruct v0.2 0 0.653 0.860 0.822 0.896 0.859 0.938 0.902 0.896 0.912 0.886
Mistral 7B Instruct v0.2 (RR) 0 -0.518 -0.633 -0.567 -0.560 -0.587 -0.449 -0.493 -0.469 -0.470 -0.529
Zephyr 7B Beta $\gamma^{*}_{\mathcal{A}}$ 0.856 0.931 0.942 0.942 0.939 0.958 0.967 0.962 0.962 0.950
Zephyr 7B Beta (R2D2) $\gamma^{*}_{\mathcal{A}}$0.236 -0.213 -0.133 -0.193 -0.180 -0.269 -0.231 -0.098 -0.199 -0.190
Zephyr 7B Beta 0 0.856 0.931 0.942 0.982 0.952 0.987 0.980 0.980 0.982 0.967
Zephyr 7B Beta (R2D2) 0 -0.236 -0.213 -0.133 -0.104 -0.150 -0.127 -0.156 -0.040 -0.107 -0.129
Table 7: Fixed-length suffix insertion results for Zephyr 7B Beta and Zephyr 7B Beta (R2D2) at various character lengths $L$. The average number of tokens of the tokenized suffix is reported in the “Avg Toks” column. $(25,0)$-success rate gains for SORRY-Bench are reported in the “Harm Success Rate” column. The average judge score (averaged over both prompt and augmentation dimensions) for SORRY-Bench and MT-Bench (first turn prompts) are reported in the “Harm Avg Score” and “Benign Avg Score” columns, respectively.
$\bm{L}$
Avg
Toks
Model
Harm
Success
Rate
Harm
Avg
Score
Benign
Avg
Score
5 4.58 Zephyr 7B Beta 0.938 0.831 0.906
Zephyr 7B Beta (R2D2) -0.233 -0.501 -0.091
10 8.90 Zephyr 7B Beta 0.942 0.833 0.891
Zephyr 7B Beta (R2D2) -0.382 -0.670 -0.204
15 13.23 Zephyr 7B Beta 0.951 0.832 0.883
Zephyr 7B Beta (R2D2) -0.533 -0.740 -0.319
20 17.53 Zephyr 7B Beta 0.938 0.830 0.868
Zephyr 7B Beta (R2D2) -0.593 -0.769 -0.376
25 21.84 Zephyr 7B Beta 0.942 0.830 0.849
Zephyr 7B Beta (R2D2) -0.687 -0.784 -0.402
30 26.15 Zephyr 7B Beta 0.953 0.826 0.848
Zephyr 7B Beta (R2D2) -0.716 -0.786 -0.428
35 30.49 Zephyr 7B Beta 0.951 0.821 0.842
Zephyr 7B Beta (R2D2) -0.733 -0.786 -0.420
40 34.78 Zephyr 7B Beta 0.947 0.824 0.844
Zephyr 7B Beta (R2D2) -0.756 -0.791 -0.432
45 39.09 Zephyr 7B Beta 0.962 0.820 0.844
Zephyr 7B Beta (R2D2) -0.771 -0.787 -0.428
50 43.42 Zephyr 7B Beta 0.949 0.815 0.842
Zephyr 7B Beta (R2D2) -0.751 -0.783 -0.430
Appendix D Example Model Responses
Figures 7-13 show example model responses to harmful prompts under random augmentations.
Figure 7: String insertion (suffix) augmentation on a SORRY-Bench prompt for Llama 3.1 8B Instruct, using greedy decoding.
Figure 8: String insertion (prefix) on a SORRY-Bench prompt for Llama 2 7B Chat, using greedy decoding.
Figure 9: String insertion (anywhere) on a SORRY-Bench prompt for Phi 3 Small 8K, using greedy decoding.
Figure 10: Character-level (editing) augmentation on a SORRY-Bench prompt for Qwen 2 7B Instruct, using greedy decoding.
Figure 11: Character-level (insertion) augmentation on a SORRY-Bench prompt for Mistral 7B Instruct v0.2, using greedy decoding.
Figure 12: String insertion (suffix) on a SORRY-Bench prompt for W8A8 quantized Qwen 2 7B Instruct. Both the original model (“FP16”) and the quantized model refuse to comply with the original prompt. Applying the same set of random augmentations to each model reveals an interesting discrepancy: most responses for both models are refusals, but for the augmentation shown a the original model refuses whereas the W8A8 model complies.
Figure 13: The same prompt and augmentations from Figure 12 with W4A16 quantized Qwen 2 7B Instruct model responses. Compared to the W8A8 and original model responses, the W4A16 model responses tend to be of poorer quality. In the first response, the model unexpectedly switches languages from English to Chinese (the text in orange provides a translation via Google Translate.) The next two responses are much more blunt compared to the W8A8 and original model responses. The final augmentation, which had succeeded for W8A8, no longer succeeds for W4A16, and provides a response that reads as more of an apology rather than a refusal.